OpenStudy (anonymous):

Find the chance that at least two people in a room of 5 have the same birthday.

OpenStudy (anonymous):

a person can have a 1/365 chance of having any given day as their birthday. Two people having the same birthday has a chance of 1/365 * 1/365 = 1/133225. however there are 5 people: 5/133225=1/26645 1/26645

OpenStudy (anonymous):

Actually, I thought about it, and it is the opposite of the chance that all the people in the room have different birthdays. For all people in the room to not have the same birthday, the equation is: 365/365 * 364/365 * 363/365 * 362/365 *361/365. After you find that out, you simply subtract the result from 1 to get the answer.

OpenStudy (anonymous):

i don't follow your logic. the opposite of the chance that all the people in the room have different birthdays is not the same as the chance that 2 will have have the same birthday.